A pyroelectric sensor has the following attributes and measurements:

Sensor capacitance = 200 μF
Sensor area = 2 mm2
Pyroelectric charge coefficient = 12 Coulombs / m2 - °C
Change in temperature across sensor faces = 6.5°C
What is the pyroelectric voltage in volts across the sensor faces?
1
Expert's answer
2020-07-13T07:30:54-0400

The charge in pyroelectric sensors depends on the pyroelectric charge coefficient, area, and temperature change:


"Q=P_QA\Delta T."

The voltage is


"V=\frac{Q}{C}=\frac{P_QA\Delta T}{C},\\\space\\
V=\frac{12(2\cdot10^{-6})6.5}{200\cdot10^{-6}}=0.78\text{ V}."